### v1

@@ 3.3.3 Theorem 3: Latency Scaling Law
$\begin{split}\tau_{\text{prefill}}&=\frac{T\cdot N\cdot\text{FLOPS}_{\text{token}}}{f_{i}}\\ \tau_{\text{decode}}&=\frac{(S-1)\cdot T\cdot N\cdot\text{FLOPS}_{\text{token}}}{f_{i}\cdot B_{i}/B_{0}}\\ \tau_{\text{io}}&=\sum_{j}\text{data\_size}_{j}/\text{BW}_{ij}\\ \tau_{\text{overhead}}&=\text{const}+\alpha\cdot\log(S)\quad\text{(heterogeneous only)}\end{split}$ (4)
Explanation: - $\tau_{\text{prefill}}$ is dominated by arithmetic operations when processing all input tokens simultaneously. This phase exhibits high arithmetic intensity (many FLOPs per byte), benefiting from high-frequency computation. The latency depends on device frequency $f_{i}$ (GHz) and the FLOP count per token (typically $2N$ for transformer inference).
@@
3.3.4 Theorem 4: Cost Scaling Law
Theorem 4.1 (Infrastructure Cost Scaling). The economic cost of inference across heterogeneous infrastructure scales as:
$\text{Cost}_{\text{total}}=\sum_{i}\left(\text{Amort}_{i}+\text{Energy}_{i}+\text{Maint}_{i}\right)$ (5)
@@ 3.3.4 Theorem 4: Cost Scaling Law
$\begin{split}\text{Amort}_{i}&=\frac{\text{HW\_Cost}_{i}}{\text{Device\_Lifetime}_{ops}}\cdot S\\ \text{Energy}_{i}&=E_{\text{total}}(S,N,T,Q,i)\cdot\text{Price}_{\text{kWh}}\\ \text{Maint}_{i}&=\text{Const}_{i}\cdot S\end{split}$ (6)
Explanation: - $\text{Amort}_{i}$ distributes hardware purchase cost across device lifetime (typically millions of operations for embedded devices). The device lifetime is device-specific: desktop CPUs have $10^{10}$ operations lifetime, embedded NPUs have $10^{13}$ operations, high-end data center GPUs have $10^{14}$ operations.
- $\text{Energy}_{i}$ multiplies total energy consumption by electricity price (typically \$0.10-\$0.15/kWh in developed countries; higher in data centers, lower in renewable-powered facilities). This directly incentivizes energy-efficient device selection.
- $\text{Maint}_{i}$ captures operational costs (cooling for GPUs, none for NPUs/CPUs in passive cooling scenarios, software licensing, etc.).
The cost per query decreases with heterogeneous placement as lower-cost devices handle suitable tasks:
$\text{Cost}_{\text{heterogeneous}}=(1-\rho)\cdot\text{Cost}_{\text{cloud}}$ (7)
where $\rho\in(0.4,0.7)$ is cost reduction factor empirically determined by optimal task-device matching. For example, a cloud query costs \$0.001 (GPU time + electricity); using heterogeneous orchestration reduces this to \$0.0003-\$0.0006 per query through optimal device selection.
3.3.5 Theorem 5: Device-Task Efficiency Compatibility
Theorem 5.1 (Hardware-Task Matching Optimality). For a task characterized by arithmetic intensity $I=\text{FLOPs}/\text{Bytes}$ and a device characterized by compute capability $C$ (FLOPS/s) and memory bandwidth $B$ (bytes/s), optimal latency is achieved when:
@@ 3.3.5 Theorem 5: Device-Task Efficiency Compatibility
Application to Inference: - Prefill with $I\approx 2N/T$ (high for large $N$ and small $T$) is compute-bound for most realistic models, benefiting from compute-optimized devices (GPUs, high-frequency CPUs). - Decode with $I\approx 1$ (memory-bound, one FLOP per byte loaded) benefits from bandwidth-optimized devices (GPUs, NPUs). - Embedding with $I\approx 0.01$ (extremely memory-bound) benefits from bandwidth and latency-optimized devices; pure CPUs are acceptable for low throughput.
This principle drives our greedy layer assignment algorithm, assigning layers based on their arithmetic intensity profile to the most suitable heterogeneous device.
@@ 3.4 Energy-Aware Task Decomposition
3.4.1 Embedding Stage
The embedding layer projects input tokens (represented as integer IDs) to continuous hidden vectors:
$\mathbf{h}_{0}=\text{Embed}(\mathbf{x})\in\mathbb{R}^{L\times d_{h}}$ (10)
where $\mathbf{x}$ is the input token sequence of length $L$, $d_{h}$ is the hidden dimension (typically 768 for BERT-scale, 4096 for GPT-3-scale), and the embedding matrix has size $V_{\text{vocab}}\times d_{h}$ (e.g., 50,000 tokens × 768 dimensions = 38.4MB for GPT-2). The operation is a simple table lookup followed by addition with positional encodings.
Characteristics: - Arithmetic intensity: Very low ($I\approx 0.01$ for $d_{h}=768$), since the compute (single addition) is minimal compared to memory access (loading embedding vectors from DRAM) - Memory access pattern: Sequential and cache-friendly for small vocabularies - FLOPs: Negligible ($\approx L$ additions for positional encoding) - Memory bandwidth requirement: $L\times d_{h}\times 2$ bytes (FP16) loaded from memory - Optimal device: CPU or NPU with fast DRAM access; unnecessary GPU overhead for such memory-bound operation
3.4.2 Transformer Decoder Layers
Each decoder block $l\in[1,L_{\text{decoder}}]$ (typically $L_{\text{decoder}}=12$ for BERT, 96 for GPT-3) performs multi-head self-attention followed by feed-forward networks:
$\begin{split}\mathbf{a}_{l}&=\text{MultiHeadAttention}(\mathbf{h}_{l-1})\\ \mathbf{h}_{l}&=\text{FFN}(\mathbf{a}_{l})+\mathbf{a}_{l}=\text{MLP}(\mathbf{a}_{l})+\mathbf{a}_{l}\end{split}$ (11)
Multi-Head Attention Mechanism:
$\text{MHA}(\mathbf{h})=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{n_{h}})\mathbf{W}^{O}$ (12)
where each attention head computes:
$\text{head}_{i}=\text{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{i}$ (13)
The query, key, and value projections are computed as:
$\mathbf{Q}_{i}=\mathbf{h}\mathbf{W}_{Q}^{i},\quad\mathbf{K}_{i}=\mathbf{h}\mathbf{W}_{K}^{i},\quad\mathbf{V}_{i}=\mathbf{h}\mathbf{W}_{V}^{i}$ (14)
where $\mathbf{W}_{Q,K,V}^{i}\in\mathbb{R}^{d_{h}\times d_{k}}$ with $d_{k}=d_{h}/n_{h}$ (attention head dimension).
Feed-Forward Network:
Following attention, a two-layer feed-forward network applies:
$\text{FFN}(\mathbf{h})=\text{GELU}(\mathbf{h}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2}$ (15)
where $\mathbf{W}_{1}\in\mathbb{R}^{d_{h}\times d_{ff}}$ (typically $d_{ff}=4d_{h}$), and $\mathbf{W}_{2}\in\mathbb{R}^{d_{ff}\times d_{h}}$. The GELU activation introduces non-linearity.
Layer Characteristics Depend on Execution Phase:
Prefill Phase (Prompt Processing): When processing all $L$ input tokens simultaneously during prefill, the sequence dimension participates in batch operations: - Attention computation: $\mathbf{Q}\mathbf{K}^{\top}$ produces an $L\times L$ matrix of attention scores, requiring $2L^{2}d_{h}$ FLOPs (matrix multiply with broadcasting) - Arithmetic intensity: $I=\frac{2L^{2}d_{h}}{L\times d_{h}\times 3}=\frac{2L}{3}$ (high for realistic $L>100$), compute-bound for $L\geq 1024$ - FFN computation: $2\times L\times d_{h}\times d_{ff}$ FLOPs (two matrix multiplies), with intensity $I\approx 2d_{ff}/3\approx 2.7$ (moderate arithmetic intensity) - Total prefill FLOPs per layer: $\approx 4Ld_{h}^{2}$ (dominated by attention quadratic term for long sequences) - Optimal device: GPU with high compute throughput (suitable for batched matrix operations)
Decode Phase (Autoregressive Generation): When generating each subsequent token autoregressively (batch dimension becomes 1): - Attention computation: $\mathbf{Q}(1\times d_{h})\times\mathbf{K}^{\top}(d_{h}\times L)$ produces logits of size $1\times L$, requiring $2Ld_{h}$ FLOPs. The KV cache stores previous tokens’ keys/values, reducing recomputation. - Arithmetic intensity: $I=\frac{2Ld_{h}}{3d_{h}}=\frac{2L}{3}$ if KV cache is in device memory, but if KV cache must be loaded from external memory, intensity drops to $I\approx 1$ (memory-bound) - FFN computation: $2d_{h}d_{ff}$ FLOPs with intensity $I\approx 2.7$ (compute-bound) - Total decode FLOPs per layer per token: $\approx 4d_{h}^{2}$ (constant, independent of sequence length) - Memory bandwidth requirement: Loading the KV cache ($2\times L\times d_{h}$ bytes) dominates for large sequence lengths - Optimal device: NPU or CPU with efficient memory hierarchy; GPU overhead diminishes as batch size = 1
3.4.3 Language Model Head and Output Projection
The language model head projects hidden states to vocabulary logits:
$\mathbf{z}=\text{LMHead}(\mathbf{h}_{L_{\text{decoder}}})=\mathbf{h}_{L_{\text{decoder}}}\mathbf{W}_{\text{out}}+\mathbf{b}_{\text{out}}$ (16)
where $\mathbf{W}_{\text{out}}\in\mathbb{R}^{d_{h}\times V_{\text{vocab}}}$ projects from hidden dimension to vocabulary size. In practice, this often shares weights with the embedding matrix (Eq. 10) to reduce model size.
Characteristics: - Arithmetic intensity: Very low ($I\approx 1$ for realistic $V_{\text{vocab}}=50k$), since the operation is a single matrix-vector multiply $(1\times d_{h})\times(d_{h}\times V_{\text{vocab}})$, FLOPs $\approx 2d_{h}V_{\text{vocab}}$, memory $\approx 2d_{h}V_{\text{vocab}}$ bytes - Memory bandwidth critical: Loading weight matrix dominates time - Optimal device: CPU or NPU with efficient memory access
@@ 3.5 Device Capability Model and Ranking
$\mathbf{d}_{i}=(M_{i}^{\text{max}},B_{i},f_{i},P_{i},n_{\text{cores},i},\lambda_{i},C_{\text{type},i},\text{priority}_{i})$ (17)
where:
  • $M_{i}^{\text{max}}$: Maximum memory capacity (GB) for storing model weights and activations
  • $B_{i}$: Memory bandwidth (GB/s), critical for memory-bound operations; examples: CPU 30GB/s, GPU 900GB/s, NPU 200GB/s
  • $f_{i}$: Compute frequency (GHz), critical for compute-bound operations; examples: CPU 3-4GHz, GPU 1-2GHz, NPU 1-2GHz
  • $P_{i}$: Peak power consumption (watts); examples: CPU 45W, GPU 300W, NPU 5-20W
  • $n_{\text{cores},i}$: Number of compute cores for FLOPS estimation
  • $\lambda_{i}$: Utilization efficiency factor (0-1), fraction of peak power used in practice for typical inference workloads
  • $C_{\text{type},i}$: Compute type classification (CPU, GPU, NPU) for architecture-specific optimizations
  • $\text{priority}_{i}$: User-specified priority (1-5), enabling preferences for power-constrained vs. latency-constrained scenarios
@@
measured in FLOPS/joule (or equivalently, joules per FLOP). For example: - Intel CPU (8 cores, 3GHz, 45W): $E_{i}=\frac{2\times 3\times 10^{9}\times 8}{45}=1.07\times 10^{9}$ FLOPS/joule - NVIDIA GPU (5120 cores, 1.5GHz, 300W): $E_{i}=\frac{2\times 1.5\times 10^{9}\times 5120}{300}=5.12\times 10^{10}$ FLOPS/joule - Intel NPU (32 cores, 2GHz, 10W): $E_{i}=\frac{2\times 2\times 10^{9}\times 32}{10}=1.28\times 10^{11}$ FLOPS/joule
Devices are sorted by efficiency (ascending joules per FLOP) for layer assignment priority, ensuring highest-efficiency devices are allocated first.
3.6 Heterogeneous Layer Assignment Algorithm
The core optimization algorithm assigns model layers to devices to minimize total energy while satisfying memory and latency constraints. The algorithm employs greedy optimization, making locally optimal choices (highest-efficiency device for each layer) that collectively yield near-optimal global allocation.
Algorithm 1 QEIL: Energy-Aware Layer Assignment
0: Model layers $\mathcal{L}$; Devices $\mathcal{D}$; Samples $S$; Tokens $T$; Latency SLA $\tau_{\max}$
0: Allocation $\mathcal{A}$ minimizing $E_{\text{total}}$
1:Step 1: Device Ranking
2:for each device $D_{i}$ do
3:  Compute efficiency $E_{i}\leftarrow\frac{2f_{i}\cdot n_{\text{cores},i}}{P_{i}}$
4:end for
5: Sort devices by efficiency ascending (most efficient first)
6:Step 2: Assign Embedding & LM Head

### v2

@@ 3.3.3 Scaling Formalism 3: Latency Scaling
$\tau_{\text{prefill}}$ $=\frac{T\cdot N\cdot\text{FLOPs}_{\text{token}}}{f_{i}}$
$\tau_{\text{decode}}$ $=\frac{(S-1)\cdot T\cdot N\cdot\text{FLOPs}_{\text{token}}}{f_{i}\cdot B_{i}/B_{0}}$
$\tau_{\text{io}}$ $=\sum_{j}\text{data\_size}_{j}/\text{BW}_{ij}$
$\tau_{\text{overhead}}$ $=\text{const}+\alpha\cdot\log(S)\quad\text{(heterogeneous only)}$ (4)
Explanation: $\tau_{\text{prefill}}$ is dominated by arithmetic operations when processing all input tokens simultaneously. This phase exhibits high arithmetic intensity (many FLOPs per byte), benefiting from high-frequency computation. The latency depends on device frequency $f_{i}$ (GHz) and the FLOP count per token (typically $2N$ for transformer inference).
@@
3.3.4 Scaling Formalism 4: Cost Scaling
Formalism 4.1 (Infrastructure Cost Scaling). We observe that the economic cost of inference across heterogeneous infrastructure scales as:
$\text{Cost}_{\text{total}}=\sum_{i}(\text{Amort}_{i}+\text{Energy}_{i}+\text{Maint}_{i})$ (5)
@@ 3.3.4 Scaling Formalism 4: Cost Scaling
$\text{Amort}_{i}$ $=\frac{\text{HW\_Cost}_{i}}{\text{Device\_Lifetime}_{\text{ops}}}\cdot S$
$\text{Energy}_{i}$ $=E_{\text{total}}(S,N,T,Q,i)\cdot\text{Price}_{\text{kWh}}$
$\text{Maint}_{i}$ $=\text{Const}_{i}\cdot S$ (6)
3.3.5 Scaling Formalism 5: Device-Task Efficiency Compatibility
Formalism 5.1 (Hardware-Task Matching Optimality). For a task characterized by arithmetic intensity $I=\text{FLOPs}/\text{Bytes}$ and a device characterized by compute capability $C$ (FLOPS/s) and memory bandwidth $B$ (bytes/s), we find that optimal latency is achieved when:
@@ 3.3.5 Scaling Formalism 5: Device-Task Efficiency Compatibility
3.4 Safety and Reliability Principles
Beyond efficiency optimization, QEIL implements a comprehensive safety and reliability framework that treats system health as a first-class constraint. Our “safety-first, capability-second” philosophy ensures that the heterogeneous orchestrator never compromises device integrity or user safety for performance gains.
3.4.1 Principle 6: Thermal Safety Constraints
Principle 6.1 (Thermal Protection). For each device $i$ with maximum safe operating temperature $T^{\max}_{i}$ and current temperature $T_{i}$, we enforce:
$T_{i}\leq\theta_{\text{throttle}}\cdot T^{\max}_{i}\quad\text{where }\theta_{\text{throttle}}=0.85$ (8)
When $T_{i}$ exceeds $\theta_{\text{throttle}}\cdot T^{\max}_{i}$, the orchestrator proactively reduces workload allocation to device $i$ by factor $(1-(T_{i}-\theta_{\text{throttle}}\cdot T^{\max}_{i})/(T^{\max}_{i}-\theta_{\text{throttle}}\cdot T^{\max}_{i}))$, redistributing work to cooler devices. This 15% safety margin prevents thermal throttling from being triggered by the hardware itself, which would cause unpredictable latency spikes, and protects against hardware damage from sustained high-temperature operation.
Implementation: Temperature monitoring is performed via device-specific interfaces: NVIDIA GPUs expose temperature through nvidia-smi and NVML; Intel CPUs/NPUs expose junction temperature through MSR registers and thermal sensors; system temperature is monitored via ACPI thermal zones. Monitoring frequency is 1Hz during normal operation and 10Hz when any device exceeds 70% of thermal limit.
3.4.2 Principle 6.2: Fault Tolerance and Graceful Degradation
Principle 6.2 (Fault-Tolerant Execution). The orchestrator maintains a device health state $H_{i}\in\{healthy,degraded,failed\}$ for each device and implements automatic recovery:
  • Failure detection: Device failures are detected through timeout monitoring (inference exceeding 10× expected latency), error rate monitoring (¿1% kernel failures over 100 inferences), and heartbeat failures (device becomes unresponsive).
  • Automatic recovery: Upon detecting device failure, the orchestrator: (1) marks the device as failedfailed, (2) redistributes pending and in-flight workloads to healthy devices within 100ms, (3) attempts device recovery (driver reset, memory clear), and (4) gradually reintroduces recovered devices starting at 50% capacity.
  • Graceful degradation: When devices fail, the system continues operating on remaining hardware with reduced throughput rather than failing entirely. The coverage-energy trade-off shifts (fewer samples per query), but the system remains responsive.
Formal guarantee: If at least one device remains healthy, QEIL guarantees continued operation with latency bounded by $\tau_{\text{degraded}}\leq\tau_{\text{optimal}}\cdot D/D_{\text{healthy}}$ where $D$ is total device count and $D_{\text{healthy}}$ is healthy device count.
3.4.3 Principle 6.3: Adversarial Robustness
Principle 6.3 (Input Validation and Anomaly Detection). To prevent adversarial inputs from causing system instability, QEIL implements defense-in-depth:
  • Input validation: Inputs are validated for: maximum sequence length (reject inputs exceeding model context window), character encoding (reject malformed UTF-8), and token rate (rate-limit to prevent resource exhaustion).
  • Output sanity checking: Model outputs are checked for: maximum generation length (hard cap at 2× expected output length), repetition detection (halt generation if >>90% token repetition over 100 tokens), and confidence anomalies (flag outputs with unusual logit distributions).
  • Resource consumption bounds: Each inference is allocated maximum memory budget ($M_{\text{max}}=1.5\times E[\text{memory}]$) and maximum time budget ($\tau_{\text{max}}=5\times E[\text{latency}]$); exceeding either triggers graceful termination.
Security note: These mechanisms protect against accidental misuse and simple adversarial attacks but are not designed to defend against sophisticated adaptive adversaries. For high-security deployments, additional hardening would be required.
@@ 3.6 Device Capability Model and Ranking
$d_{i}=(M^{\max}_{i},B_{i},f_{i},P_{i},n_{\text{cores},i},\lambda_{i},C_{\text{type},i},\mathbf{T^{\max}_{i}},\text{priority}_{i})$ (10)
@@
Note on thermal limits: $T^{\max}_{i}$ is obtained from device specifications and represents the maximum junction temperature before hardware damage occurs. We enforce operation at 85% of this limit to provide safety margin.
3.7 Optimization Problem Formulation with Hardware and Safety Constraints
The complete QEIL optimization synthesizes all objectives with hardware-specific constraints derived from our experimental platform and safety requirements:
$\min_{A}\quad$ $E_{\text{total}}(A)=\sum_{i}(E_{\text{prefill},i}+E_{\text{decode},i})$
s.t. $\sum_{l\in L_{i}}M_{l}\leq M^{\max}_{i}$
$M_{\text{CPU}}\leq 127\text{ GB},\quad M_{\text{NPU}}\leq 20\text{ GB}$
$M_{\text{GPU1}}\leq 96.2\text{ GB},\quad M_{\text{GPU2}}\leq 72.7\text{ GB}$
$B_{\text{CPU}}=100\text{ GB/s},\quad B_{\text{NPU}}=50\text{ GB/s}$
$P_{\text{CPU}}\leq 45\text{ W},\quad P_{\text{GPU}}\leq 300\text{ W},\quad P_{\text{NPU}}\leq 25\text{ W}$
$\tau_{\text{total}}(A)\leq\tau_{\max}$
$C(S,N,T)\geq C_{\min}$
$\textbf{T}_{i}(A)\leq 0.85\cdot T^{\max}_{i}\quad\forall i\quad\text{(thermal safety)}$ (12)
These constraints ensure feasibility on our experimental edge platform: Intel Core Ultra 9 285HX processor (8 cores, 2.80 GHz), 128 GB system RAM (127 GB usable), Intel AI Boost NPU with 20 GB dedicated storage, NVIDIA RTX PRO 5000 Blackwell GPU (96.2 GB total VRAM), and Intel Graphics GPU (72.7 GB shared memory). Memory, power, and thermal limits reflect realistic edge deployment scenarios where devices operate under strict resource budgets while maintaining inference quality, throughput, and safe operation.
Greedy Algorithm Justification: We use a greedy layer assignment algorithm rather than optimal methods (ILP, dynamic programming) for three reasons: (1) Computational efficiency: Greedy assignment runs in $O(L\cdot D)$ time where $L$ is layer count and $D$ is device count, enabling real-time reallocation; ILP formulations are NP-hard. (2) Near-optimality: For our workloads, the greedy solution achieves within 5% of the ILP optimum (validated on subset of experiments), as layer energy costs are approximately uniform and assignment has submodular structure. (3) Safety compatibility: Greedy assignment naturally accommodates dynamic safety constraints (thermal throttling, device failures) through re-execution, whereas optimal solutions would require expensive re-computation.
4 Ablation Studies
To validate the robustness of QEIL’s design choices and quantify the contribution of individual components, we conduct comprehensive ablation studies across seven dimensions: (1) scaling exponent stability test, (2) controlled heterogeneity ablation isolating the effect of hardware diversity, (3) component contribution analysis, (4) variance and reproducibility assessment, (5) energy and latency breakdown analysis with real-time orchestrator visualization, (6) cross-dataset robustness evaluation, and (7) safety and reliability validation. These ablations follow best practices on scaling relationships and systems optimization Brown et al. (2024); Hoffmann et al. (2022).
4.1 Scaling Exponent Stability Test ($\beta$ Stability)
A critical assumption in QEIL is that the coverage scaling exponents $\beta_{N}$ and $\beta_{S}$ (approximately 0.7) are stable across transformer model families. We validate this by measuring $\beta$ independently for each model and analyzing sensitivity to hyperparameter variations.
Table 1: Scaling Exponent $\beta$ Stability Across Model Families. Values computed via nonlinear least-squares fitting of $C(S)=1-\exp(-\alpha S^{\beta})$ across $S\in\{1,5,10,15,20\}$ samples. 95% confidence intervals computed via bootstrap resampling (1000 iterations). Note: Overlapping confidence intervals indicate we cannot statistically distinguish model-specific exponents; we interpret this as evidence of approximate stability rather than exact equality.
Model $\beta$ (fitted) 95% CI $R^{2}$
GPT-2 (125M) 0.68 [0.64, 0.72] 0.994
Granite-350M 0.71 [0.67, 0.75] 0.991
Qwen2-0.5B 0.69 [0.65, 0.73] 0.993
Llama-3.2-1B 0.72 [0.68, 0.76] 0.996
LFM2-2.6B 0.70 [0.66, 0.74] 0.995
Mean 0.70 [0.66, 0.74] 0.994
Key Finding: The scaling exponent $\beta$ exhibits consistent values across all tested transformer families, with mean $\beta=0.70\pm 0.02$ and all confidence intervals overlapping. Statistical interpretation: The overlapping confidence intervals mean we cannot reject the null hypothesis that all models share the same $\beta$ value. We interpret this as evidence that $\beta$ is approximately architecture-invariant within the transformer family on our tested hardware, while acknowledging that larger sample sizes might reveal model-specific differences. The high $R^{2}$ values (>> 0.99) confirm excellent fit quality, suggesting the power-law relationship captures the empirical data well.
Sensitivity to Sample Range: We additionally tested whether $\beta$ varies when computed over different sample ranges:
Table 2: Scaling Exponent Sensitivity to Sample Budget Range.
Sample Range $\beta$ (GPT-2) $\beta$ (Llama) $\Delta\beta$
$S\in[1,10]$ 0.66 0.70 0.04
$S\in[1,20]$ 0.68 0.72 0.04
$S\in[5,50]$ 0.71 0.74 0.03
$S\in[10,100]$ 0.73 0.76 0.03
The exponent shows mild increase (+0.05+0.05) when computed over larger sample ranges, consistent with diminishing returns at higher sample budgets. Importantly, the cross-model difference $\Delta\beta$ remains small ($\leq 0.04$), confirming architecture-invariance within our tested range.
4.2 Controlled Heterogeneity Ablation
To isolate the contribution of heterogeneous orchestration from confounding factors (e.g., simply using better hardware), we conduct a controlled ablation comparing three configurations using identical models and workloads:
Table 3: Controlled Heterogeneity Ablation: Isolating the Effect of Hardware Diversity. All configurations use GPT-2 (125M) with $S=20$ samples on WikiText-103. “Homogeneous GPU” runs all operations on NVIDIA RTX PRO 5000; “Homogeneous NPU” runs all operations on Intel AI Boost NPU; “Heterogeneous (QEIL)” uses our orchestration strategy. Energy and latency measured via hardware counters; coverage via pass@k evaluation.
Configuration Pass@k (%) Energy (kJ) Latency (ms) IPW Power (W) PPP
Homogeneous GPU 59.5 43.1 1.73 0.149 402.5 16.85
Homogeneous NPU 58.2 31.8 2.41 0.312 186.4 14.21
Homogeneous CPU 57.8 38.6 3.12 0.187 309.2 12.94
Heterogeneous (QEIL) 70.0 22.5 1.34 0.718 83.5 20.74
$\Delta$ vs. Best Homogeneous +10.5pp -29.2% -22.5% +130% -55.2% +23.1%
Critical Insight: The heterogeneous configuration outperforms all homogeneous baselines across every metric simultaneously. This rules out the alternative hypothesis that QEIL’s gains stem solely from using better hardware—rather, the gains arise from intelligent task-device matching that exploits complementary hardware strengths:
  • Coverage gains (+10.5pp): Heterogeneous orchestration enables more effective sample diversity by reducing per-sample latency variance through device-specialized execution paths.
  • Energy reduction (-29.2% vs. NPU): Even compared to the most energy-efficient homogeneous device (NPU), heterogeneous execution achieves lower total energy by routing compute-bound prefill to GPU (higher throughput) and memory-bound decode to NPU (lower power).
  • Latency improvement (-22.5% vs. GPU): Despite using lower-power devices for some operations, heterogeneous orchestration achieves lower latency than GPU-only execution by avoiding GPU memory bandwidth bottlenecks during autoregressive decode.
4.3 Component Contribution Analysis
We isolate the contribution of each QEIL component by progressively enabling features:
Table 4: Component Contribution Analysis: Incremental Effect of QEIL Features on GPT-2 (125M).
Configuration Pass@k (%) Energy (kJ) IPW
Baseline (GPU-only) 59.5 43.1 0.149
+ Device Ranking 61.2 38.7 0.178
+ Prefill/Decode Split 65.8 29.4 0.412
+ Greedy Layer Assignment 68.3 25.1 0.584
+ Adaptive Sample Budget 69.2 23.4 0.672
+ Safety Constraints 70.0 22.5 0.718
Findings:
  • Device ranking provides modest gains (+1.7pp coverage, -10.2% energy) by prioritizing efficient devices.
  • Prefill/decode disaggregation is the largest single contributor (+4.6pp coverage, -24.0% energy), validating our phase-aware task decomposition.
  • Greedy layer assignment provides further optimization (+2.5pp, -14.6% energy) through fine-grained allocation.
  • Adaptive sample budget enables coverage gains (+0.9pp, -6.8% energy) by dynamically adjusting samples under energy constraints.
  • Safety constraints provide additional gains (+0.8pp, -3.8% energy) by preventing thermal throttling and ensuring stable operation—demonstrating that safety-first design improves rather than degrades performance by avoiding unpredictable thermal-induced slowdowns.
4.4 Variance and Reproducibility Analysis
To assess result stability, we report variance across multiple experimental runs:
Table 5: Variance Analysis: Standard Deviation Across 10 Independent Runs for GPT-2 (125M) with QEIL Energy-Aware Configuration.
Metric Mean Std Dev CV (%)
Pass@k (%) 70.0 0.82 1.17
Energy (kJ) 22.5 0.41 1.82
Latency (ms) 1.34 0.03 2.24
IPW 0.718 0.015 2.09
Power (W) 83.5 1.24 1.49
All metrics exhibit low coefficient of variation (CV << 2.5%), confirming high reproducibility. The variance in coverage (CV = 1.17%) is particularly important as it validates that the pass@k improvements are statistically significant rather than noise artifacts. Note: This variance analysis is performed on a single hardware platform; cross-platform variance would likely be higher and requires future investigation.
4.5 Cross-Model Consistency Validation
We verify that ablation findings generalize across model families:
Table 6: Cross-Model Ablation Consistency: Heterogeneous vs. Best Homogeneous Baseline Across All Models.